Stock Y has a beta of 1.4 and an expected return of 13 percent. Stock Z has a beta of 0.85 and an expected return of 10.4 percent. Required: What would the risk-free rate have to be for the two stocks to be correctly priced relative to each other? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16).) Risk-free rate % Suggestions: We need to set the reward-to-risk ratios of the two assets equal to each other, which is:.......(equation needed) We then cross multiply to get: Solving for the risk-free rate, we find:
Answer #1

We need to set the reward-to-risk ratios of the two assets equal to each other, which is:

(0.13 – Rf) / 1.4 = (0.104 – Rf) / 0.85

We can cross multiply to get:

0.85(0.13 – Rf) = 1.4(0.104 – Rf)

Solving for the risk-free rate, we find:

0.1105 – 0.85Rf = 0.1456 – 1.4Rf

Rf = 0.0638, or 6.38%